# Unveiling the Landscape of LLM Deployment in the Wild: An Empirical Study

### v1

@@ IV-B2 Functional Coverage of Exposed Endpoints
The functionality exposed through open API endpoints reveals not only the intended capabilities of LLM frameworks but also the extent to which internal operations are externally accessible, whether intentionally or not. As shown in Table V, the degree of exposure varies substantially across both frameworks and functionality categories. Some frameworks expose broad functionality: for instance, Jan and llama.cpp each respond to over 10 distinct endpoints across text generation, embedding, file operations, and fine-tuning. ComfyUI returns valid responses for 11 categories, including session and queue management, reflecting its interactive, stateful design. In contrast, frameworks like OpenWebUI and Text Generation WebUI expose few usable endpoints despite implementing many, suggesting stricter access controls or incomplete external integration. Across frameworks, the most commonly exposed functionality is text and chat generation, observed in 11 out of 13 frameworks. Model operations (e.g., listing and loading models) are also frequently reachable in 8 frameworks. However, sensitive features like fine-tuning, moderation, and knowledge base querying remain rare; for instance, only Jan and vLLM expose fine-tuning endpoints, and OpenWebUI supports RAG but with no responsive endpoints. The observed exposure patterns reveal trade-offs between usability and security. Frameworks intended for local or development use often leak internal APIs by default without authentication or isolation, while production-oriented systems show tighter surface control. Even endpoints related to queue status or system configuration, while not critical, can still cause unintended exposure of internal system states.
@@ IV-B3 Per-Endpoint Result Representation
All API responses are first normalized into a unified schema capturing deployment framework, endpoint category, endpoint path, response type, and potential security relevance. This abstraction enables consistent analysis across frameworks with differing designs and naming conventions. Based on this schema, we analyze Ollama as a representative case. Table VI summarizes endpoint responsiveness across 384 observed invocations. Text and chat completion endpoints responded in 57.29% of cases. While often active, their partial availability suggests exposure conditioned on runtime state or configuration, potentially leaking system behavior to unauthenticated clients. The embedding endpoint was accessible in only 2.60% of attempts. Though rarely enabled, its occasional exposure still poses risk, as embedding vectors may support inference attacks even in limited contexts. Model management endpoints showed uneven availability. Listing local models was frequently possible (75.26%), likely due to its role in coordination. Pull and push operations were moderately available (66.67%, 54.17%), while destructive actions like delete, copy, and create responded in fewer than 30% of cases, suggesting partial lockdowns, though enforcement remains inconsistent. System-level endpoints, including version and runtime model queries, responded in under 3% of cases. This likely reflects hardening efforts, but their occasional exposure highlights the need for stricter endpoint visibility controls. The results highlight that in Ollama deployments, many endpoints, including those that expose model metadata, are accessible without authentication, allowing unauthorized users to retrieve sensitive information and revealing a critical lack of access control.
TABLE VI: Responsiveness of Ollama API Endpoints.
@@
Text/Chat Gen /api/generate Text Completion 220 57.29%
/api/chat Chat Completion 220 57.29%
Embedding /api/embeddings Generate Embedding 10 2.60%
Model Ops /api/tags List Local Models 289 75.26%
/api/pull Pull Model 256 66.67%
/api/push Push Model 208 54.17%
/api/delete Delete Model 109 28.39%
/api/copy Copy Model 80 20.83%
/api/show Show Model Info 36 9.38%
/api/create Create Model 33 8.59%
/api/running List Loaded Models 9 2.34%
System Config /api/version Get Ollama Version 9 2.34%
  • Note: Based on 384 sampled LLM service invocations.
Takeaway 2: Our findings reveal widespread security weaknesses in real-world LLM deployments. A large portion of services, such as Ollama and Llamafile, allow over 80% of unauthenticated API requests, exposing critical operations like text generation and model management (e.g., 75.26% model listing success). Even when frameworks appear restrictive, inconsistencies and partial exposures persist. This indicates that current deployment practices lack systematic access control and minimal exposure principles, leaving many services vulnerable to unauthorized access, information leakage, and misuse.
[…]
Text Gen WebUI 0.31% 0.31% 0.31%
GPT4All 2.40%
Llamacpp 4.53% 2.27%
Llamafile 16.57%
Jupyter Notebook 13.19%
  • 1
[…]
@@
Model Information Disclosure /embeddings 7 1.87%
/history 25 6.67%
System Configuration Disclosure /system_stats 17 4.53%
Unauthorized & Resource Abuse /queue 4 1.07%
/prompt 12 3.20%
Vulnerabilities & Reverse /object_info 6 1.60%
/extensions 35 9.33%
Sensitive Content Generation /prompt 12 3.20%
@@
    Based on analysis of 375 sampled ComfyUI endpoints. # Cases and Case Rates indicate confirmed security risks..
IV-C Security and Risk Analysis
As shown in Table VII, this section systematically analyzes potential security risks in LLM deployment frameworks. Each subcategory combines observed cases with deeper security insights to reveal not only surface-level exposures but also underlying attack vectors and systemic vulnerabilities.
IV-C1 Model Information Disclosure
Model information disclosure is among the most prevalent security risks across LLM deployment frameworks, as indicated in Table VII. For example, the /show endpoint can reveal detailed model metadata, while the /history endpoint may leak prior inference workflows, outputs, and fine-tuning traces, compromising user privacy and exposing proprietary data. Notably, in Ollama, the /api/tags endpoint exhibits an exceptionally high exposure rate of 70.57%, highlighting the severity of this risk even in widely adopted deployment platforms. The /embeddings endpoint (1.87% exposure) in ComfyUI further illustrates this problem by exposing the list of loaded textual inversion embeddings, such as Bad-Hands-XL, inadvertently revealing deployment purposes. Similarly, in llama.cpp, the /v1/models endpoint may disclose detailed model information, including the actual deployment paths of .gguf model files, significantly amplifying the risk of targeted attacks. Such leakage not only facilitates customized attacks but also heightens the risks of model exploitation, prompt injection, and unauthorized access. In severe cases, attackers could reconstruct user intents, manipulate outputs, or compromise the deployment’s integrity and confidentiality.
@@ IV-C2 Hardware and System Configuration Disclosure
Disclosure of system configuration details further enlarges the attack surface by providing adversaries with valuable intelligence about the underlying environment. ComfyUI’s /system_stats endpoint (4.53%) reveals operating system types, total and available memory, GPU specifications (e.g., RTX 4090 with 24GB VRAM), and ComfyUI version information. Such granular system insights enable attackers to craft hardware-specific resource exhaustion attacks, such as GPU memory flooding, or identify platform-specific vulnerabilities (e.g., Windows NT kernel exploits). Alarmingly, among instances exposing /system_stats, 41.28% were found to be publicly accessible via --listen 0.0.0.0 without any authentication mechanisms. When combined with the presence of high-performance GPUs, such exposures create ideal conditions for unauthorized resource exploitation, including GPU hijacking, cryptocurrency mining, large-scale model inference, or persistent backdoor implantation.
@@ IV-C3 Unauthorized Access and Resource Abuse
Unauthorized access and resource abuse represent direct threats to the availability and stability of deployed services. Endpoints such as /queue (1.07%) and /prompt (3.20%) allow unauthenticated users to submit inference tasks or monitor system load. The ability to observe task queues enables attackers to perform real-time load sensing, identifying idle periods when resource-draining or prompt injection attacks can be launched with minimal resistance. Furthermore, submitting crafted prompt graphs to /prompt without authentication exacerbates the risk: attackers could overload the system with computationally expensive tasks, rapidly depleting GPU memory and causing service outages (denial-of-service attacks). By continuously monitoring queue states, attackers can also infer operational patterns over time, improving the precision and persistence of their abuse strategies. In the absence of fine-grained access control and rate-limiting, such systems remain highly vulnerable to long-term degradation and operational hijacking.
@@ IV-C4 Vulnerabilities and Reverse Engineering
The exposure of internal modules and metadata creates direct opportunities for vulnerability discovery and reverse engineering. In platforms like ComfyUI, endpoints such as /extensions (9.33%) and /object_info (1.60%) reveal detailed information about installed nodes, system structure, and behaviors. With such insights, attackers can reconstruct internal workflows, pinpoint critical components like PythonEvalNode, and exploit weaknesses such as insecure APIs or insufficient validation. Nodes handling external communication (e.g., GeminiAPINode) may leak API keys, while backend modules (e.g., cm-api.js) expose surfaces for command injection. Similar risks arise in another framework. Jupyter Notebook instances exposing the /api endpoint disclose version data (e.g., “5.5.0”), allowing attackers to link targets to known vulnerabilities, including unauthorized API access, unauthenticated WebSocket RCE, and token bypass. Fingerprinting deployments and mapping them to public exploits significantly accelerates attack development. Passive metadata collection can quickly escalate into full system compromise, underscoring the critical risks of seemingly minor exposures.
@@ IV-C5 Sensitive Content Generation
Takeaway 3: Security risk analysis shows that model information disclosure, system leaks, unauthorized access, vulnerabilities, and sensitive content generation are prevalent across LLM deployment frameworks, though unevenly. Notably, frameworks like ComfyUI expose endpoints across all major risk categories, indicating broad and systemic weaknesses. The persistence of such exposures indicates that insecure deployment practices are widespread, and securing LLM systems demands rethinking default configurations, strengthening access controls, and minimizing exposure surfaces beyond patching individual flaws.
@@ V Discussion
V-A Root Causes of Interface Exposure
The widespread exposure of LLM deployment interfaces stems from a combination of insecure defaults, weak boundary enforcement, and operational misunderstandings. Many frameworks are designed with the assumption of a trusted local environment, where binding to “localhost” is seen as sufficient. However, in practice, deployments often occur in containerized or multi-user settings, where services are inadvertently exposed to broader networks without authentication or access control. This risk is amplified by a usability-first design culture. To simplify prototyping, many tools expose powerful endpoints by default and provide minimal guidance on securing them. Documentation often emphasizes quick interaction through HTTP APIs but overlooks the security implications of commands related to model loading, prompt submission, or file manipulation. When authentication options exist, they are frequently disabled by default or poorly documented. Another key factor is the insufficient separation between user-level interactions and system-level control. Endpoints that allow prompt execution, plugin loading, or system inspection are often directly accessible without input validation or role-based restrictions. In extensible platforms, dynamically loaded modules operate without sandboxing, creating opportunities for privilege escalation or internal reconnaissance. Finally, interface exposure is exacerbated by limited user awareness. Developers and deployers often assume that local services are inherently secure, ignoring risks introduced by container networking, port forwarding, or LAN visibility. Without monitoring or audit mechanisms, these oversights can lead to persistent vulnerabilities and external abuse.
V-B Recommendations and Best Practices
Improving the security posture of local LLM frameworks requires coordinated efforts across tooling, usage, and community practices.
Tool developers: enforce secure defaults and restrict exposure. Developers of LLM-serving frameworks should adopt secure-by-default principles as a baseline. All interfaces should bind to the local loopback address unless explicitly configured otherwise, and sensitive endpoints involving model management, file access, or system inspection should require authentication or confirmation even in local testing environments. Debugging interfaces and metadata endpoints should remain disabled by default and only be exposed through deliberate configuration. In extensible platforms that support third-party modules or scripts, execution should be sandboxed, and plugin behavior must be scoped and controlled to prevent arbitrary access or privilege escalation. Documentation must emphasize the presence of security-critical endpoints, provide clear guidance for enabling authentication and authorization, and promote best practices during local or production deployment. Security features should be surfaced rather than buried as optional configurations.
System deployers: treat local interfaces as externally reachable. Operators deploying LLM systems must treat every interface, even those running locally, as potentially accessible to untrusted users, especially in shared-host, containerized, or local network environments. Interfaces should be explicitly reviewed and minimized, unnecessary endpoints disabled, and authentication mechanisms activated wherever available. Networking configurations must be handled with caution. Container networking, port forwarding, and cross-device accessibility can expose services far beyond what is intended. Interfaces that handle prompt submission, queue monitoring, or tool invocation must be closely audited and protected against misuse or overuse. Regular security testing should be incorporated into deployment pipelines, including automated scans for exposed interfaces and simulated jailbreak attempts. Logging mechanisms should be comprehensive enough to support post-incident analysis and real-time anomaly detection.
Community stakeholders: promote secure deployment norms. As LLM frameworks transition rapidly from research prototypes to production-grade systems, the surrounding ecosystem must keep pace by establishing a culture of secure deployment. Insecure defaults in upstream tools often propagate downstream without scrutiny, increasing the systemic risk of widespread vulnerabilities. The community should actively develop and maintain secure deployment templates, configuration checklists, and audit frameworks tailored to LLM-serving use cases. Public repositories should annotate models and frameworks with deployment guidance, security metadata, and interface descriptions to help users apply safe configurations. More broadly, deployment security should be treated on par with performance and usability, supported by transparency, disclosure, and community-led testing.
V-C Limitation

### v2

@@ IV-B2 Functional Coverage of Exposed Endpoints
The functionality exposed through open API endpoints reveals not only the intended capabilities of LLM frameworks but also the extent to which internal operations are externally accessible, whether intentionally or not. As shown in Table V, the degree of exposure varies substantially across both frameworks and functionality categories. Some frameworks expose broad functionality: for instance, Jan and llama.cpp each respond to over 10 distinct endpoints across text generation, embedding, file operations, and fine-tuning. ComfyUI returns valid responses for 11 categories, including session and queue management, reflecting its interactive, stateful design. In contrast, frameworks like OpenWebUI and Text Generation WebUI expose few usable endpoints despite implementing many, suggesting stricter access controls or incomplete external integration. Across frameworks, the most commonly exposed functionality is text and chat generation, observed in 11 out of 13 frameworks. Model operations (e.g., listing and loading models) are also frequently reachable in 8 frameworks. However, sensitive features like fine-tuning, moderation, and knowledge base querying remain rare; for instance, only Jan and vLLM expose fine-tuning endpoints, and OpenWebUI supports RAG but with no responsive endpoints. Exposure patterns highlight the trade-off between usability and security. Frameworks designed for local or development use often default to exposing internal APIs, lacking authentication or isolation, while production systems enforce stricter controls. Even non-critical endpoints, such as those for queue status or configuration, can inadvertently leak internal state.
@@ IV-B3 Per-Endpoint Result Representation
All API responses are normalized into a unified schema encompassing deployment framework, endpoint category, endpoint path, response type, and potential security relevance, enabling consistent analysis across different frameworks. Based on this schema, we analyze Ollama as a representative case study. Table VII summarizes the endpoint responses for 155,424 observed calls. Text and chat completion endpoints responded in only 0.19% of cases, indicating limited exposure based on runtime state or configuration, but still potentially revealing system behavior. The embedding endpoint was accessible in 0.22% of attempts; while rarely exposed, its exposure still poses a risk because embeddings can be used in inference attacks. Model management endpoints showed uneven availability.Show model infomation and list running models were most frequently possible (9.94%, 14.40%), while listing local models was moderately possible (4.50%), likely due to its role in coordination. Pull and push operations were low available (0.20%, 0.19%), while destructive actions like delete, copy, and create responded in fewer than 2% of cases, suggesting partial lockdowns, though enforcement remains inconsistent. System-level endpoints, including version and runtime model queries, responded in 14.48% of cases, the same level to list running models. This likely reflects hardening efforts, but their occasional exposure highlights the need for stricter endpoint visibility controls. The results highlight that in Ollama deployments, many endpoints are accessible without authentication, allowing unauthorized users to retrieve sensitive information and revealing a critical lack of access control.
TABLE VI: Risk Categorization (Min $\sim$ Max Percentage) of Different LLM Deployment Frameworks.
ComfyUI 5.26% $\sim$ 10.54% 5.64% 5.06% $\sim$ 5.26% 5.35% $\sim$ 10.09% 5.06%
Ollama 0.00% $\sim$ 14.48% 0.19% $\sim$ 1.28% 0.19% $\sim$ 0.22%
Text Gen WebUI 0.29% 0.29% 0.29%
GPT4All 2.87% 5.10%
Llamacpp 3.36% $\sim$ 6.99% 6.71% 11.92%
Llamafile 8.98% $\sim$ 10.19% 9.47%
Jupyter Notebook 8.06% 0.16%
  • 1
    Each cell shows the observed minimum $\sim$ maximum percentage of API endpoints posing the corresponding risk category. If minimum equals maximum, only a single value is shown. “–” indicates no relevant risk observed. Detailed case rates for ComfyUI endpoints are summarized in Table VIII.
Answer to RQ2: API exposure and access control vary across LLM frameworks. Ollama and Llamafile responded to over 10% of unauthenticated requests, while other frameworks, such as Open WebUI and Gradio, responded at rates below 0.1%. Most frameworks exposed basic text generation functionality, but some also leaked model or system information, particularly Ollama, where this rate exceeded 14%. Many frameworks relied on permissive defaults, exposing deployments to the risk of unauthorized access.
TABLE VII: Responsiveness of Ollama API Endpoints.
@@
Text/Chat Gen /api/generate Text Completion 299 0.19%
/api/chat Chat Completion 300 0.19%
Embedding /api/embeddings Generate Embedding 343 0.22%
Model Ops /api/tags List Local Models 7,001 4.50%
/api/pull Pull Model 312 0.20%
/api/push Push Model 299 0.19%
/api/delete Delete Model 306 0.20%
/api/copy Copy Model 1986 1.28%
/api/show Show Model Info 15,452 9.94%
/api/create Create Model 311 0.20%
/api/ps List Running Models 22,375 14.40%
/api/running List Loaded Models 0 0.00%
System Config /api/version Get Ollama Version 22,503 14.48%
  • Note: Based on 155,424 sampled LLM service invocations.
IV-C RQ3: Security and Risk Analysis
As shown in Table VI, this section systematically analyzes potential security risks in LLM deployment frameworks. Each subcategory combines observed cases with deeper security insights to reveal not only surface-level exposures but also underlying attack vectors and systemic vulnerabilities.
IV-C1 Model Information Disclosure
Model information disclosure is among the most prevalent security risks across LLM deployment frameworks, as indicated in Table VI. For example, the /show endpoint can reveal detailed model metadata, while the /history endpoint may leak prior inference workflows, outputs, and fine-tuning traces, compromising user privacy and exposing proprietary data. Notably, in Ollama, the /api/show endpoint exhibits an exceptionally high exposure rate of 14.40%, highlighting the severity of this risk even in widely adopted deployment platforms. The /embeddings endpoint (5.26% exposure) in ComfyUI further illustrates this problem by exposing the list of loaded textual inversion embeddings, such as Bad-Hands-XL, inadvertently revealing deployment purposes. Similarly, in llama.cpp, the /v1/models endpoint may disclose detailed model information, including the actual deployment paths of .gguf model files, significantly amplifying the risk of targeted attacks. Such leakage not only facilitates customized attacks but also heightens the risks of model exploitation, prompt injection, and unauthorized access. In severe cases, attackers could reconstruct user intents, manipulate outputs, or compromise the deployment’s integrity and confidentiality.
@@ IV-C1 Model Information Disclosure
Model Information Disclosure /embeddings 330 5.29%
/history 658 10.54%
System Configuration Disclosure /system_stats 352 5.64%
Unauthorized & Resource Abuse /queue 316 5.26%
/prompt 328 5.06%
Vulnerabilities & Reverse /object_info 630 10.09%
/extensions 334 5.35%
Sensitive Content Generation /prompt 328 5.06%
@@
    Based on analysis of 6,241 sampled ComfyUI endpoints. # Cases and Case Rates indicate confirmed security risks.
@@ IV-C2 Hardware and System Configuration Disclosure
Disclosure of system configuration details further enlarges the attack surface by providing adversaries with valuable intelligence about the underlying environment. ComfyUI’s /system_stats endpoint (5.64%) reveals operating system types, total and available memory, GPU specifications (e.g., RTX 4090 with 24GB VRAM), and ComfyUI version information. Such granular system insights enable attackers to craft hardware-specific resource exhaustion attacks, such as GPU memory flooding, or identify platform-specific vulnerabilities (e.g., Windows NT kernel exploits). Alarmingly, among instances exposing /system_stats, 41.28% were found to be publicly accessible via --listen 0.0.0.0 without any authentication mechanisms. When combined with the presence of high-performance GPUs, such exposures create ideal conditions for unauthorized resource exploitation, including GPU hijacking, cryptocurrency mining, large-scale model inference, or persistent backdoor implantation.
@@ IV-C3 Unauthorized Access and Resource Abuse
Unauthorized access and resource abuse represent direct threats to the availability and stability of deployed services. Endpoints such as /queue (5.26%) and /prompt (5.06%) allow unauthenticated users to submit inference tasks or monitor system load. The ability to observe task queues enables attackers to perform real-time load sensing, identifying idle periods when resource-draining or prompt injection attacks can be launched with minimal resistance. Furthermore, submitting crafted prompt graphs to /prompt without authentication exacerbates the risk: attackers could overload the system with computationally expensive tasks, rapidly depleting GPU memory and causing service outages (denial-of-service attacks). By continuously monitoring queue states, attackers can also infer operational patterns over time, improving the precision and persistence of their abuse strategies. In the absence of fine-grained access control and rate-limiting, such systems remain highly vulnerable to long-term degradation and operational hijacking.
@@ IV-C4 Vulnerabilities and Reverse Engineering
The exposure of internal modules and metadata creates direct opportunities for vulnerability discovery and reverse engineering. In platforms like ComfyUI, endpoints such as /extensions (5.35%) and /object_info (10.09%) reveal detailed information about installed nodes, system structure, and behaviors. With such insights, attackers can reconstruct internal workflows, pinpoint critical components like PythonEvalNode, and exploit weaknesses such as insecure APIs or insufficient validation. Nodes handling external communication (e.g., GeminiAPINode) may leak API keys, while backend modules (e.g., cm-api.js) expose surfaces for command injection. Similar risks arise in another framework. Jupyter Notebook instances exposing the /api endpoint disclose version data (e.g., “5.5.0”), allowing attackers to link targets to known vulnerabilities, including unauthorized API access, unauthenticated WebSocket RCE, and token bypass. Fingerprinting deployments and mapping them to public exploits significantly accelerates attack development. Passive metadata collection can quickly escalate into full system compromise, underscoring the critical risks of seemingly minor exposures.
@@ IV-C5 Sensitive Content Generation
Answer to RQ3: Security risk analysis shows that model information disclosure, system leaks, unauthorized access, vulnerabilities, and sensitive content generation are prevalent across LLM deployment frameworks, though unevenly. Notably, frameworks like ComfyUI expose endpoints across all major risk categories, indicating broad and systemic weaknesses. The persistence of such exposures indicates that insecure deployment practices are widespread, and securing LLM systems demands rethinking default configurations, strengthening access controls, and minimizing exposure surfaces beyond patching individual flaws.
@@ V Discussion
V-A Recommendations and Best Practices
Improving the security posture of local LLM frameworks requires coordinated efforts across tooling, usage, and community practices. For tool developers. Frameworks should adopt a secure-by-default approach. APIs should bind to localhost by default, require explicit configuration for external access, and issue warnings when sensitive endpoints (e.g., model deletion, file upload) are accessed without authentication. Interface metadata, such as OpenAPI schemas, can define access roles and scopes, enabling future integration with RBAC systems. Optional security features should be documented clearly and made easier to configure. For end users. Users should avoid assuming that “local” tools are secure by default. Exposing endpoints, including those within containers or LAN environments, should be done with caution. Recommended practices include restricting interfaces, enabling authentication where available, and disabling unnecessary endpoints. Users should consult community guides and treat LLM tools as potentially network-accessible services. For the broader ecosystem. The blurred line between research prototypes and production deployments poses systemic risks. Insecure defaults in upstream tools can propagate into downstream projects, widening the attack surface. Community resources such as deployment checklists, secure-by-default templates, and best practices can help align tool design with real-world deployment needs. Security should be integral to LLM system design, not treated as an afterthought.
V-B Threats to Validity